OpenStudy (anonymous):

Find the measure of each interior angle of a regular dodecagon A 30 B 180 C150 D 360

OpenStudy (anonymous):

sum of interior angles=(n-2)*180..n=number of sides

OpenStudy (anonymous):

dodecagon has 12 sides...

OpenStudy (anonymous):

sum of angles=(12-2)*180=1800

OpenStudy (anonymous):

so one angle will have 1800/12=150 degree

OpenStudy (anonymous):

@dokolo54

OpenStudy (anonymous):

thanks